OpenStudy (anonymous):

Solving this equation x=4y + 1

OpenStudy (anonymous):

For what?? I did not get you..

OpenStudy (anonymous):

I need to figure out how 4y + 1 = 4y-1 have the same solution. I left out a portion of the question.

OpenStudy (anonymous):

Wait, what is your original question??

OpenStudy (anonymous):

Do the equation x=4y +1 and x = 4y-1, have the same solution

OpenStudy (anonymous):

They should not be..

OpenStudy (anonymous):

I am thinking not, too. However, I need to explain why there is no solution.

OpenStudy (anonymous):

The thing is that: x = 4y + 1 x = 4y - 1 Now subtract the two equations: 0 = 0 + 2 Here you have no x and y when you subtract. No Solution just means that the two given lines equations are not intersecting each other. hence no solution and at the same time they are not parallel you must remember this..

OpenStudy (anonymous):

That is what I left out is to remember they do not intersect with each other. Thank you...

OpenStudy (anonymous):

are you doing elimination?

OpenStudy (anonymous):

No Solution means they are not intersecting and are not parallel or coinciding lines.. Infinitely many solutions means that they are parallel and are overlapping lines..

OpenStudy (anonymous):

Okay, Thank you writing this down to remember.

OpenStudy (anonymous):

You are welcome dear..

OpenStudy (anonymous):

To Iforgot, I need to explain why there is no solution for the both equations. thank you for your help!

OpenStudy (anonymous):

o okies